OpenStudy (anonymous):

Use FOIL to find the product. (2x^5 -9)(x^8 -1)=

OpenStudy (mathteacher1729):

FOIL First Outside Inside Last First = 2x^5 times x^8 = 2x^13 Outside = 2x^5 times -1 = -2x^5 Inside = -9 times x^8 = -9x^8 Last = -9 times -8 = 72 Now add up all the terms: 2x^13 - 2x^5 - 9x^8 + 72 There are no "like terms" (same powers of x) so we are done. :)